OpenStudy (anonymous):

Which circles have an area of ft2? Choose all answers that are correct. A. a circle with a radius of 6 ft B. a circle with a diameter of 6 ft C. a circle with a radius of 12 ft D. a circle with a diameter of 12 ft

OpenStudy (anonymous):

@soprano.h.d0816

OpenStudy (xmissalycatx):

Do you know the area formula for a circle?

OpenStudy (misssmartiez):

\[3.14r^2 \]

OpenStudy (misssmartiez):

That is the given formula to find the area of a circle. So, what do you suppose is the answer? (remember, darling, r^2 isn't r x 2. It is r x r.

OpenStudy (mathmale):

Kindly provide the missing (numerical) information: "Which circles have an area of ft2?"

OpenStudy (anonymous):

@peak

OpenStudy (anonymous):

none of them r correct bby

OpenStudy (anonymous):

i checked 3 times. none of them equal 2ft

OpenStudy (anonymous):

wat no

OpenStudy (anonymous):

Which circles have an area of ft2? Choose all answers that are correct. A. a circle with a radius of 6 ft B. a circle with a diameter of 6 ft C. a circle with a radius of 12 ft D. a circle with a diameter of 12 ft

OpenStudy (anonymous):

\[36\]

OpenStudy (anonymous):

pi

OpenStudy (anonymous):

please help

OpenStudy (mathstudent55):

The formula for the area of a circle is \(A = \pi r^2\) You are told the area is \(36 \pi\) Set \(\pi r^2 = 36 \pi\) Divide both sides by \pi. What do you get?

OpenStudy (mathstudent55):